Calculate number of moles of hydrogen gas contained in 20 L flask at 27°C exerting pressure 70 cm Hg. (R = 0.082 L atm K-1mol-1)


1
Expert's answer
2020-08-05T05:05:53-0400

pv = nRT

n = (p * v) / (R * T) = (20*0.0921) / (0.082*300) = 0.075 mol.

Answer: 0.075 mol.